OpenStudy (anonymous):

-2(x + 3) = -2(x + 1) - 4

OpenStudy (anonymous):

so you distribute -2x-6=-2x-2-4 -2x-6=-2x-6 since the equations are equal x can equal all real numbers

OpenStudy (anonymous):

okay thank you! can you help me with another?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

One month, a music site observed that 60% of the people who downloaded songs from its site downloaded Q Sam's latest single. The equation below represents this information, where x represents the total number of people who downloaded songs from the site that month. x = 0.6x + 384 How many people who downloaded songs from the site that month downloaded Q Sam's latest single? 576 240 144 960

OpenStudy (anonymous):

wrong one sorry

OpenStudy (anonymous):

hhaaha i chose the wrong one hold on

OpenStudy (anonymous):

Solve for x: 3 over 4 x + 5 over 8 = 4x x = __________________ Write your answer as a fraction in simplest form. Use the "/" symbol for the fraction bar.

OpenStudy (anonymous):

so the equation is \[\frac{ \frac{ 3 }{ 4 }x +5}{ 8 }=4x\] right?

OpenStudy (anonymous):

no its 3/4X+5/8=4x

OpenStudy (haseeb96):

this question doesnot exist???

OpenStudy (anonymous):

?

OpenStudy (anonymous):

ok the so you subtract 3/4x from both sides to get 13/4x=5 multiply by 4/13 on both sides x=20/13

OpenStudy (anonymous):